What mass of KX (Mm= 150 g/mol) is required to prepare 331 mL of a pH = 12.68 solution? The pKa of HX is 11.75

g KX
Answer #1

mass of KX = 22.6 g

Explanation

pKa of HX = 11.75

pKb of X- = 14 - pKa of HX

pKb of X- = 14 - 11.75

pKb of X- = 2.25

Kb = 10-pKb

Kb = 10-2.25

Kb = 5.62 x 10-3

pH = 12.68

pOH = 14 - pH

pOH = 14 - 12.68

pOH = 1.32

[OH-] = 10-pOH

[OH-] = 10-1.32

[OH-] = 0.04786 M

initial concentration of KX = ([OH-]2 / Kb) + [OH-]

initial concentration of KX = [(0.04786 M)2 / 5.62 x 10-3] + (0.04786 M)

initial concentration of KX = 0.4555 M

volume of solution = 331 mL = 0.331 L

moles of KX = (initial concentration of KX) * (volume of solution in Liter)

moles of KX = (0.4555 M) * (0.331 L)

moles of KX = 0.151 mol

mass of KX = (moles of KX) * (molar mass KX)

mass of KX = (0.151 mol) * (150 g/mol)

mass of KX = 22.6 g